a body in a straight line travels 2 m in the first two seconds and 2.2m in the next four seconds with constant retardation. what will be its velocity at the end of the seventh second from the start?

Grade:9

1 Answers

Sunil Kumar FP
9 years ago
let the initial velocity be u
velocity at the end of second sth second=v1
velocity at the end of 6thsecond=v2
velocity at the end of the 7thsecond=v3
retardation=@
v1=u-2@
v2=u-6@
v3=u-7@
now
2=(u-2@)2-1/2@4
u-@=1
2.2=(u-2@)4-1/2@16
@=3/20
now v3=u-7@
=0.1m/s
thanks and regards
sunil kr
askIITian faculty